OpenStudy (anonymous):

How do you set up this? you are planning to start a small business that will require an investment of $90,000. YOU have found some people who are willing to share equally in the venture. if you found three more people, each person's share will decrease by $2500, How many people have found so far?

OpenStudy (oaktree):

Okay. So let's say that you've found n people. So the share of each person is \[\frac{ $90,000 }{ n }\]Following so far?

OpenStudy (anonymous):

yes

OpenStudy (oaktree):

Great. So then if we find three more people, the number of people is n+3, so the shares are\[\frac{ $90,000 }{ n+3 }\]So we have the equation \[\frac{ 90,000 }{ n } = \frac{ 90,000 }{ n+3 } + 2500\]Good?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got it

OpenStudy (oaktree):

Good. So you have the answer?

OpenStudy (anonymous):

yes thanks so much for your help!!

OpenStudy (oaktree):

No problem.